# Orca: Browsing at Scale Through User-Driven and AI-Facilitated Orchestration Across Malleable Webpages

### v1

@@ Abstract.
Web-based activities are fundamentally distributed across webpages. However, conventional browsers with stacks of tabs fail to support operating and synthesizing large volumes of information across pages. While recent AI systems enable fully automated web browsing and information synthesis, they often diminish user agency and hinder contextual understanding. Therefore, we explore how AI could instead augment users’ interactions with content across webpages and mitigate cognitive and manual efforts. Through literature on information tasks and web browsing challenges, and an iterative design process, we present a rich set of novel interactions with our prototype web browser, Orca. Leveraging AI, Orca supports user-driven exploration, operation, organization, and synthesis of web content at scale. To enable browsing at scale, webpages are treated as malleable materials that humans and AI can collaboratively manipulate and compose into a malleable, dynamic, and browser-level workspace. Our evaluation revealed an increased “appetite” for information foraging, enhanced user control, and more flexibility in sensemaking across a broader information landscape on the web.
@@
ccs: Human-centered computing Human computer interaction (HCI)ccs: Information systems Information systems applicationsccs: Information systems Browsers
@@ 1. Introduction
Recent advances in Large Language Models (LLMs) have brought transformative changes to how we interact with information on the web. For example, OpenAI’s Operator integrates a chat interface with a classic browser, which helps translate users’ natural language prompts into a sequence of operations on a browser (ope, 2025). Systems like Deep Research further remove the user from the browsing and synthesis process, delivering a final report (dee, 2025). While these approaches can reduce the cognitive and manual effort of browsing, managing, and synthesizing content, direct user interaction with web content remains irreplaceable for several reasons. First, compared to AI-generated summaries, directly engaging with content allows users to consume information in its original context, enabling a more accurate understanding of the problem space (Pirolli and Card, 1995). Second, user-driven browsing lets users stay in control of information sources, steer their evolving exploration directions, preserve a sense of agency, and leave room for serendipity (Foster and Ford, 2003; Teevan et al., 2004; Suh et al., 2023; Pirolli and Card, 2005).
Therefore, instead of using AI to take the users away from the web, we aim to explore enhancing user-driven browsing by preserving user engagement with web content and employing AI to mitigate undesired cognitive and manual effort. We explore how browsers can be redesigned to enable browsing at scale. Our key approach is to treat webpages as “malleable materials”—elements that can be viewed in parallel and flexibly extracted—and the browser as a “malleable space”–an environment where these materials can be dynamically reorganized and composed into a custom workspace. With AI facilitation, users can then view, organize, and manipulate content across webpages in a scalable and flexible manner, which we define as “orchestration.”
@@
To ground our design exploration that enables browsing at scale, we synthesized the literature that identified the merits of manual effort and cognitive costs during web browsing and sensemaking. Reflecting on the activities involved—from starting to synthesizing—we break down the broader goal of browsing-at-scale into specific interaction challenges and opportunities.
Building on these insights, we present a rich set of novel interactions resulting from our exploration with Orca  [^1] Resources are available at https://hci.ucsd.edu/orca and https://jiang.pl/orca., a prototype web browser. Orca supports user-driven exploration, operation, organization, and synthesis of web content at scale while leveraging AI to help users stay in control. For example, with Orca, users can open a large number of links in batches to broaden the exploration, with autonomous agents concurrently deployed across tabs, to accelerate repetitive interactions occurring across webpages. Webpages can be transformed into more succinct forms and flexibly organized into different structures based on their content and users’ context. Finally, information across pages can be synthesized on demand, supporting progressive sensemaking for the dynamic webpages and exploration. As a result, Orca helps users build and manage a dynamic information space, with the web as its foundation.
@@ 2. Problem Framing
This work emerges amid rapid advancements in LLMs and their applications. Fully automated information agents like OpenAI Deep Research, Perplexity, and Manus now demonstrate considerable capabilities to forage information, synthesize, and generate comprehensive reports from simple user queries, transforming how we access the information (Venkit et al., 2024; dee, 2025; per, 2022; man, 2025).
However, HCI and sensemaking research have long emphasized the benefits of the information foraging and synthesizing process itself (Rieh et al., 2016; Teevan et al., 2004). Therefore, this work takes a different stance—We do not want to replace browsing with fully automated agents. Instead, we explore leveraging AI’s capabilities and integrating them into each activity in the information-seeking and sensemaking process, to reduce cognitive and interaction costs while preserving the cognitive benefits of direct information engagement.
@@
2.1. Starting
Informational tasks often start without well-defined goals and domain-specific knowledge and language that are needed to translate users’ vague needs into concrete queries (Aula and Russell, 2008; Marchionini, 2006; Hearst, 2009; Ellis, 1989). As a result, selecting starting points and crafting initial queries can be demanding and often leads to confusion and ineffective paths, requiring users to iteratively adjust their queries and sources throughout the process (Aula and Russell, 2008; Bates, 1989; White and Roth, 2009; Kuhlthau, 1991). On the other hand, being able to flexibly restart and reformulate queries allows users to adaptively anchor and reorient their exploration and gain a deeper understanding of their goal and necessary vocabulary (Vakkari, 2001; Rieh et al., 2014).
@@ 2.4. Synthesizing and Transacting
Meanwhile, web and UI automation agents powered by LLMs show increasing promise in reducing manual efforts in web browsing (Zhou et al., 2023; Müller and Žunič, 2024; He et al., 2024; Koh et al., 2024; Deng et al., 2023; Huq et al., 2025). These agents, including OpenAI Operator and Claude Computer Use, translate natural language instructions into concrete actions (like clicks and scrolling), continuously analyze unseen interfaces, and adaptively determine the next steps to take without requiring extensive end-user training, scripting, or step-by-step monitoring (ope, 2025; com, 2024). While their performance in the wild still needs substantial improvement, the rapid advancement of LLMs suggests we are stepping into a promising future where automation agents may reliably and swiftly navigate convoluted web environments, gather information, and complete complex tasks.
@@ 3. Summary and Design Goal
Informational tasks on the web involve fluidly interleaved activities from starting and scouting, to synthesizing and transacting. We synthesize key insights from prior literature in Table 1, outlining the typical tasks, user goals, associated costs and benefits, and potential opportunities for facilitation.
Prior research has highlighted the value of manual effort in these activities: It supports anchored exploration aligned with evolving and personal needs (White and Roth, 2009; Ellis, 1989; Teevan et al., 2004); enables serendipitous discovery (Foster and Ford, 2003); and fosters progressive understanding of the problem space (Teevan et al., 2004; Suh et al., 2023; Pirolli and Card, 2005). Manual interaction also contributes to building reusable knowledge, understanding the reasoning behind the final results, and maintaining user agency and trust (Teevan et al., 2004; Pirolli and Card, 1995).
Meanwhile, web-based information tasks require users to repeatedly navigate through distributed webpages and manage large volumes of fragmented information and context. Traditional tabbed browsing, however, prioritizes working within each individual page and barely helps with cross-page activities, amplifying navigation and synthesizing costs (Chang et al., 2021a; Ma et al., 2023; Brown and Shillner, 1995).
Therefore, this work aims to facilitate a user-driven approach to the web information space to preserve the merits of manual effort, while reducing cognitive and interaction costs carried by the distributed nature of web-based tasks.
@@ 3.1. Design Goal: Browsing at Scale
We aim to enable browsing at scale to reduce repetitive low-level operations and constant context switching, supporting tasks across many webpages. Browsing at scale means:
@@
    Instead of restarting the search with new queries and sources repeatedly, or navigating into a single thread or subpage at a time, users could start and navigate at scale by expanding from their current context to surface a broad set of relevant resources and exploring multiple pages in parallel.
@@
    Instead of making sense of information from fragmented sources manually, users could synthesize at scale with summaries that integrate content of multiple sources.
@@ 4. Approach
We explore a two-fold design approach to facilitate at-scale, user-driven browsing, as detailed in this section.
4.1. Webpages as Malleable Materials, and Browser as a Malleable Space
We address the limitations of traditional tabbed browsing, which treats each page as an isolated, primary workspace. Tabs are organized in a linear stack, making it difficult to view or operate across multiple pages at once. Additionally, individual pages often contain unrelated content, requiring users to repeatedly sift through noise and perform the same navigation and interaction steps.
Therefore, we reimagine webpages as “malleable materials”—elements that can be viewed at varying scales, transformed, and extracted. Correspondingly, we reconceptualize the browser as a “malleable space”–an environment where these materials can be flexibly composed into custom workspaces. Our approach unfolds across these two dimensions as follows:
Malleable webpage materials: parallel view and customized extraction. To enable at-scale actions across pages, malleable webpages support being viewed at a “zoomed-out” scale, allowing users to monitor and interact with many pages simultaneously without losing context (Section 5.1). To reduce distraction and support focused scanning, content across pages can be selectively extracted—surfacing only the relevant information in a customized view aligned with users’ tasks and informational needs (Section 5.4).
Malleable browser space: dynamic organization and composition. Moving beyond the constraints of linear tabs, malleable webpages can be flexibly reorganized, grouped, and composed into various structures, such as tables. These groupings and structures allow users to distinguish webpages for different tasks, visualize relationships across content, and adapt layouts to fit evolving workflows—forming a malleable, fluid, and task-centric browser environment from previously disconnected webpages (Section 5.2).
4.2. User-Driven and AI-Facilitated Orchestration
On the other hand, we aim to enable browsing at scale through a user-driven approach that preserves cognitive benefits and maintains user control over evolving goals and preferences. We explore interface designs and interaction techniques that empower users to navigate, operate on, and synthesize across multiple malleable webpages simultaneously—a process we define as “orchestration.” In this process, AI plays a facilitating role: it accelerates reach, reduces repetitive manual effort, and adapts content to aid sensemaking, while allowing the user to steer and control the collaborative orchestration across malleable pages. Specifically, we explore three collaborative models between the user and AI:
AI scales up, user scales down. To help users quickly bootstrap and expand their exploration across sources, AI interprets the context and content of the pages being browsed and “scales up” the exploration by identifying relevant directions and alternative entry points, and extending the context of available information. The user then “scales down” by reviewing the available materials, filtering based on relevance or their interest, and actively steering the exploration to align with their evolving goals and needs. This collaborative model accelerates the discovery of a broadened information landscape without removing users from the loop, keeping them in control of the direction and depth of exploration (Section 5.3).
AI surfaces, user deep-dives. To support progressive sensemaking that aligns with users’ evolving inquiries, AI helps distill content across malleable webpages into succinct forms and synthesizes cross-source summaries that surface key points in context. Users remain in control of interpretation by selectively deep-diving into areas of interest, posing follow-up questions, and iteratively bringing in new or alternative content for updated extraction and synthesis. This collaborative model supports efficient scanning while preserving direct engagement with source material, enabling a personalized and evolving sensemaking process grounded in intermediate, contextualized digests (Sections 5.45.5).
AI executes, user monitors. To reduce the cognitive and manual overhead of repetitive interactions, AI carries out low-level UI operations in parallel. Users remain in charge by monitoring, guiding, and intervening when needed. They can evaluate results, adjust direction, and override automation in real time. This collaborative model allows users to delegate and expedite tedious tasks while maintaining oversight and control (Section 5.6).
@@ 5. Browsing at Scale with Orca
We present our design exploration to enable browsing at scale with our prototype browser, Orca (Figure 2).
@@ 5.1. View at Scale: Spatial Overview Across Multiple Webpages
With traditional tabbed browsing, users struggle to constantly jump between multiple linearly stacked tabs to track different browsing threads. To address this, Orca supports showing multiple webpages in parallel by displaying them as “shapes” on the Web Canvas (Figure 2.b). Users can view pages simultaneously and rearrange them freely by dragging them around—much like using a graphics editor such as Figma (fig, 2016). For example, users might cluster pages into separate groups to keep different tasks apart, then cross-reference information or merge groups to work across tasks easily when needed. Unlike prior browsers that organize abstracted webpages spatially (using favicons or titles), Orca keeps webpages fully interactive to support live, cross-page interactions, going beyond earlier spatial tools used for code or writing (Ayers and Stasko, 1995; Bragdon et al., 2010; Jiang et al., 2023a; Suh et al., 2023; Bederson and Hollan, 1994; Agarawala and Balakrishnan, 2006; Henderson and Card, 1986; Wright et al., 2006).
Common graphics editing interaction techniques, such as range selection, also help users reposition multiple pages at once or apply batch operations across selected pages (Masson et al., 2025). Utilizing the spatial layout, users can drag any link from a page onto the canvas to open it at a specific location, while clicking the link opens it next to the original page for easy side-by-side viewing. Orca displays the actual content of multiple pages at once, not just titles or favicons. This helps users quickly get an overview and easily refind pages by recognizing visual cues from the content itself. Users can also zoom in or out, or pan the canvas, to adjust the viewport and focus on different webpages of interest.
As a result, users compose a dynamic, customized workspace at the browser level using multiple webpages. For example, when finding a hotel for a trip, instead of switching back and forth between tabs, users can include multiple pages in view to scan through the differences, move discarded options aside, combine with flight booking webpages and cross-check matches, and zoom in on selected ones to confirm the details.

### v2

@@ Abstract.
Web-based activities span multiple webpages. However, conventional browsers with stacks of tabs cannot support operating and synthesizing large volumes of information across pages. While recent AI systems enable fully automated web browsing and information synthesis, they often diminish user agency and hinder contextual understanding. We explore how AI could instead augment user interactions with content across webpages and mitigate cognitive and manual efforts. Through literature on information tasks and web browsing challenges, and an iterative design process, we present novel interactions with our prototype web browser, Orca. Leveraging AI, Orca supports user-driven exploration, operation, organization, and synthesis of web content at scale. To enable browsing at scale, webpages are treated as malleable materials that humans and AI can collaboratively manipulate and compose into a malleable, dynamic, and browser-level workspace. Our evaluation revealed an increased “appetite” for information foraging, enhanced control, and more flexible sensemaking across a broader web information landscape.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790335isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Human computer interaction (HCI)$c$s: Information systems Information systems applicationsccs: Information systems Browsers
@@
A diagram illustrates the features of the Orca browser in a five-step workflow. Step 1, viewing, shows thumbnails of several different websites. Step 2, navigating, shows a grid of four web pages. Step 3, organizing, depicts the grid view transforming into a table. Step 4, extracting, shows highlighted data elements within the pages in the grid. Step 5, operating and synthesizing, shows colorful cursors pointing at different pages with a command bar, representing an operation across multiple sources.
@@ 1. Introduction
Recent advances in Large Language Models (LLMs) have brought transformative changes to how we interact with information on the web. For example, OpenAI’s Operator integrates a chat interface with a classic browser, which helps translate users’ natural language prompts into a sequence of operations on a browser (40). Systems like Deep Research further remove the user from the browsing and synthesis process, delivering a final report (39). While these approaches can reduce the cognitive and manual effort of browsing, managing, and synthesizing content, direct user interaction with web content remains irreplaceable for several reasons. First, compared to AI-generated summaries, directly engaging with content allows users to consume information in its original context, enabling a more accurate understanding of the problem space (Pirolli and Card, 1995). Second, user-driven browsing lets users stay in control of information sources, steer their evolving exploration directions, preserve a sense of agency, and leave room for serendipity (Foster and Ford, 2003; Teevan et al., 2004; Suh et al., 2023; Pirolli and Card, 2005).
Therefore, instead of using AI to take users away from the web, we aim to enhance web-based information foraging and synthesis—the most common and demanding uses of browsers (Pirolli and Card, 1995; Sellen et al., 2002; Chang et al., 2021a)—by preserving direct engagement with web content and applying AI to reduce manual and cognitive burden. We explore how web browsers can support browsing at scale for such tasks, allowing users to interact with many webpages in parallel rather than one at a time.
Our key approach builds on prior work on canvas-based interactions and malleable software environments by treating webpages as “malleable materials”—elements that can be viewed in parallel at different scales, summarized, and flexibly extracted—and the browser as a “malleable space”—a browser-level workspace where these materials can be viewed simultaneously, spatially arranged, grouped, and composed into task-centric structures like grids, and summarized on demand (Klokmose et al., 2015; Bederson and Hollan, 1994; Wright et al., 2006; Litt and Jackson, 2020). By breaking the walls of individual webpages, this malleable workspace lets users keep all relevant pages for a task in view while flexibly reorganizing them as goals evolve. AI facilitation further enables “orchestration,” where users organize, manipulate, and synthesize content across webpages with the help of AI agents in a scalable, flexible manner.
@@
To ground our design exploration, we synthesize literature that identifies both the merits of manual effort and its cognitive costs in web browsing and sensemaking for information foraging and synthesis tasks (Pirolli and Card, 1995, 2005). Reflecting on the activities involved—from starting to synthesizing—we break down browsing-at-scale into specific interaction challenges and opportunities.
We prototype the Orca browser to instantiates this malleable workspace. Instead of isolated tabs, Orca places pages side by side on a canvas, where users can open large batches of links to broaden exploration, deploy autonomous agents across tabs to accelerate repetitive interactions, transform webpages into succinct representations, flexibly organize them into structures based on content and context, and synthesize information across pages on demand to support progressive sensemaking. Orca leverages AI facilitation and interactive features to help users build and manage a dynamic information space, with the web as its foundation.
@@ 2. Problem Framing
This work emerges amid rapid advancements in LLMs and their applications. Fully automated deep research agents like OpenAI Deep Research, Perplexity, and Manus now forage, synthesize, and generate comprehensive reports from simple user queries, transforming how we access information (P. N. Venkit, P. Laban, Y. Zhou, Y. Mao, and C. Wu (2024); 39; 82; 67).
However, HCI and sensemaking research have long emphasized the benefits of the information foraging and synthesizing process (Rieh et al., 2016; Teevan et al., 2004). Therefore, this work takes a different stance—We do not want to replace browsing with fully automated agents. Instead, we explore leveraging AI’s capabilities and integrating them into each activity in the information-seeking and sensemaking process, to reduce cognitive and interaction costs while preserving the cognitive benefits of direct information engagement.
@@
2.1. Starting
Informational tasks often start without well-defined goals and domain-specific knowledge and language that are needed to translate users’ vague needs into concrete queries (Aula and Russell, 2008; Marchionini, 2006; Hearst, 2009; Ellis, 1989). As a result, selecting starting points and crafting initial queries can be demanding and often lead to confusion and ineffective paths, requiring users to iteratively adjust their queries and sources throughout the process (Aula and Russell, 2008; Bates, 1989; White and Roth, 2009; Kuhlthau, 1991). On the other hand, being able to flexibly restart and reformulate queries allows users to adaptively anchor and reorient their exploration and gain a deeper understanding of their goal and necessary vocabulary (Vakkari, 2001; Rieh et al., 2014).
@@ 2.4. Synthesizing and Transacting
Meanwhile, web and UI automation agents powered by LLMs show increasing promise in reducing manual efforts in web browsing (Zhou et al., 2023; Müller and Žunič, 2024; He et al., 2024; Koh et al., 2024; Deng et al., 2023; Huq et al., 2025). These agents, including OpenAI Operator and Claude Computer Use, translate natural language instructions into concrete actions (like clicks and scrolling), continuously analyze unseen interfaces, and adaptively determine the next steps to take without requiring extensive end-user training, scripting, or step-by-step monitoring (40; 38). While their performance in the wild still needs substantial improvement, the rapid advancement of LLMs suggests we are stepping into a promising future where automation agents may reliably and swiftly navigate convoluted web environments, gather information, and complete complex tasks.
@@
2.5. Related Work
We briefly review other system work that our research builds on.
2.5.1. Canvas- and spatial-based interfaces
Our work builds on a long line of systems that use spatial canvases and multi-workspace environments to help people manage complex information. Early systems such as Rooms, Pad++, and BumpTop explored multiple virtual workspaces, zoomable canvases, and physics-inspired desktops to cope with limited screen space and window thrashing (Henderson and Card, 1986; Bederson and Hollan, 1994; Agarawala and Balakrishnan, 2006). MosaicG visualized web history as a two-dimensional map of visited pages to support revisitation and orientation (Ayers and Stasko, 1995). Code Bubbles and other working-set interfaces showed how collections of lightweight fragments arranged on a canvas can improve code understanding and reduce navigation overhead (Bragdon et al., 2010). Tools such as Figma demonstrate how shared canvases for many artifacts support parallel, non-linear work in everyday practice (31). Recent LLM-based systems such as Graphologue and Sensecape further leverage spatial layouts and multilevel structures to externalize, reorganize, and explore model-generated content (Jiang et al., 2023a; Suh et al., 2023).
These efforts demonstrate the value of spatial canvases for organizing many items and supporting non-linear exploration. In this work, we explore embedding webpages on a canvas and coupling spatial arrangement with AI-facilitated operations across them, enabling parallel automation and synthesis at the browser level.
2.5.2. Malleable software and web interfaces
Malleable software research has long envisioned systems that users can adapt to their own tasks, data, and workflows (Kay and Goldberg, 1977). Haystack proposed unified data models and customizable interface layers to integrate heterogeneous information and support personalized information management (Adar et al., 1999). Smalltalk enabled users to inspect and modify the behavior and appearance of GUI objects, and Webstrates treated web documents as shared, dynamic media that can be restructured and recomposed across clients (Klokmose et al., 2015; Kay, 1993). More recently, systems such as Jelly explore activity-based, data-model-driven customization, where users can reshape information interfaces through natural language and direct manipulation as their tasks evolve (Cao et al., 2025). Work on malleable web interfaces further argues for treating websites as adaptable surfaces that can be reshaped to better fit user goals and cross-site workflows (Min et al., 2025; Litt and Jackson, 2020).
In this work, we treat webpages as malleable materials in a dynamic workspace for the users to forage and make sense of web information at scale. AI agents extract, synthesize, and operate across many sites in parallel to enable users’ interaction with underlying webpages and orchestration through the malleable workspace.
@@ 3. Summary and Design Goal
Web information tasks involve fluidly interleaved activities from starting and scouting, to synthesizing and transacting across webpages. We synthesize key insights from prior literature in Table 1, outlining typical tasks, user goals, associated costs and benefits, and opportunities for facilitation in this information foraging and sensemaking process.
Prior research has highlighted the value of manual effort in these activities: it supports anchored exploration aligned with evolving and personal needs (White and Roth, 2009; Ellis, 1989; Teevan et al., 2004); enables serendipitous discovery (Foster and Ford, 2003); and fosters progressive understanding of the problem space (Teevan et al., 2004; Suh et al., 2023; Pirolli and Card, 2005). Manual interaction also contributes to building reusable knowledge, understanding the reasoning behind final results, and maintaining user agency and trust (Teevan et al., 2004; Pirolli and Card, 1995).
Meanwhile, web-based tasks require users to repeatedly navigate through distributed webpages and manage large volumes of fragmented information and context. Traditional tabbed browsing, however, prioritizes working within each individual page and barely supports cross-page activities, amplifying navigation and synthesis costs (Chang et al., 2021a; Ma et al., 2023; Brown and Shillner, 1995).
Therefore, this work aims to facilitate a user-driven approach to the web information space that preserves the merits of manual effort while reducing the cognitive and interaction costs imposed by the distributed nature of web-based tasks. We target information-foraging and sensemaking workflows that span tens of webpages, and seek to support them through a malleable browser-level workspace for browsing at scale.
@@
A screenshot of the Orca browser’s two-part user interface. The interface is divided into a narrow, foldable side panel on the left and a large Web Canvas on the right. The side panel shows a website and a row of tabs at its bottom; one tab is highlighted in purple to indicate it is currently selected. The Web Canvas is a freeform canvas where multiple webpage windows are spatially arranged and grouped.
@@ 3.1. Design Goal: Browsing at Scale
We aim to enable browsing at scale to reduce repetitive low-level operations and constant context switching in cross-page tasks. In our work, browsing at scale means that users can treat a large set of related pages as a single, malleable workspace and:
@@
    Instead of repeatedly restarting search with new queries or drilling into one thread or subpage at a time, users can start and navigate at scale by expanding from their current context to surface many relevant resources and explore multiple pages in parallel.
@@
    Instead of making sense of information from fragmented places manually, users could synthesize at scale with summaries that integrate content of multiple sources.
@@ 4. Approach
We explore a design approach that supports browsing at scale by (1) reconceptualizing webpages as “malleable materials” and the browser as a “malleable space” that together form a task-centric workspace (Klokmose et al., 2015), and (2) employing AI as a facilitator that scales users’ reach, extraction, and operation across that workspace while keeping them in control as an “orchestrator.”
First, we address the limitations of traditional tabbed browsing, which treats each page as an isolated primary workspace organized in a linear stack, making it difficult to view or operate across multiple pages at once and forcing users to repeatedly sift through unrelated content. We explore treating webpages as malleable materials that can be viewed at varying scales in parallel, selectively extracted into compact representations, and reorganized into structures such as grids, stacks, and tables (Sections 5.15.2). This malleable browser space lets users arrange, group, and recombine pages into custom workspaces aligned with their tasks, so that cross-page comparison, filtering, and sensemaking happen in one shared context rather than across scattered tabs.
Second, we support user-driven, AI-facilitated orchestration across this malleable workspace. Instead of end-to-end automation that removes users from the loop, AI helps scale up exploration by expanding from the current context to related sources and alternatives (Section 5.3), surface structure by extracting fields and synthesizing summaries across pages to aid sensemaking (Sections 5.45.5), and execute repetitive low-level operations in parallel across selected pages (Section 5.6). Users remain responsible for scaling down—filtering, organizing, deep-diving, and steering automation—so that browsing at scale emerges as a collaborative process between user and AI over malleable webpages, rather than as fully delegated web automation.
@@ 5. Browsing at Scale with Orca
We present Orca, a browser designed to support user-driven web information foraging at scale (Figure 2).
To illustrate how Orca streamlines complex browsing workflows, we follow Leela, a user planning a family trip to Paris. Through her journey, we describe how Orca transforms the isolated webpages into a malleable workspace for gathering, organizing, and synthesizing information.
@@ 5.1. View at Scale: Spatial Overview Across Multiple Webpages
Two webpages are shown side-by-side in the Orca interface. Each webpage has its own set of controls beneath it, including back and forward navigation buttons and a URL bar. Above the pages is a shared contextual menu with icons for actions that apply to both selected pages, such as pinning them or arranging them in a grid. A cursor is shown pressing the ‘Pin to the sidebar’ button.
Leela plans a family trip to Paris. She searches for hotels and drags candidates onto the Web Canvas. With Orca, she places family-friendly options at the top, groups budget ones on the left, and moves backups into a corner. She zooms out to see all options, then zooms in to read a few reviews in detail. She later sees that some hotels do not meet her criteria—she range-selects them and closes them all at once.
Traditional tabbed browsing fragments context and content across a linear sequence of pages, making it difficult to maintain context when web information branches and interweaves. To address this, Orca supports showing multiple webpages in parallel by displaying them as “shapes” on the Web Canvas (Figure 2.b).
Being able to easily build and maintain a view of many pages side by side helps users understand, track, and synthesize information at scale. As a result, users zoom, pan, and lay out webpages in ways that reflect their ongoing tasks within an ad hoc, customized workspace that better fits their ongoing task, for example, grouping hotels by price, flights by stops, or papers by topic. While some content may appear smaller and harder to read or operate on, Orca supports extracting, operating on, and synthesizing this content with AI facilitation, which we introduce in the following subsections. This allows users to utilize Web Canvas to forage and orchestrate at a higher level.
On the other hand, because Orca shows the actual content of each page rather than only titles or thumbnails, users rely on rich visual cues for refinding and comparison. Range selection and batch actions enable users to reposition or act on many pages at once (Masson et al., 2025; Kay and Goldberg, 1977). Dragging a link from a page onto the canvas opens it at a chosen position, while clicking opens it adjacent to the source page for immediate side-by-side inspection (Badam et al., 2019; Beaudouin-Lafon et al., 2012).